# Enhancing Oscillator-Based Ising Machine Models with Amplitude Dynamics and Polynomial Interactions

### v1

@@ Abstract
We present an oscillator model with both phase and amplitude dynamics for oscillator-based Ising machines that addresses combinatorial optimization problems with polynomial cost functions of arbitrary order. Our approach addresses fundamental limitations of previous oscillator-based Ising machines through a mathematically rigorous formulation with a well-defined energy function and corresponding dynamics. The model demonstrates monotonic energy decrease and reliable convergence to low-energy states. Empirical evaluations on 3-SAT problems show significant performance improvements over existing phase-amplitude models. Additionally, we propose a flexible, generalizable framework for designing higher-order oscillator interactions, from which we derive a practical method for oscillator binarization without compromising performance. This work strengthens both the theoretical foundation and practical applicability of oscillator-based Ising machines for complex optimization problems.
@@ I INTRODUCTION
The limitations of traditional von Neumann architectures in efficiently tackling combinatorial optimization problems have spurred the exploration of alternative computing paradigms. Among these, Ising machines, inspired by the Ising model from statistical mechanics, have emerged as a promising approach. The versatility of Ising machines extends across a diverse range of applications, including portfolio optimization [1][2], drug discovery [3], vehicle routing [4, 5, 6], fraud detection [7], clustering [8], power grid optimization [9], wireless communication [10, 11], image restoration [12] and machine learning [13][14][15][16]. By mapping the decision variables of an optimization problem onto interacting spins [17], Ising machines can efficiently search for solutions by allowing the system to naturally evolve towards its lowest energy state. This process corresponds to finding a configuration of spins that minimizes the system’s energy, which directly translates to a high-quality solution to the original optimization problem. This approach offers the potential to significantly accelerate the solution of complex optimization problems that are intractable for classical computers.
@@
Ising machines based on networks implementing the Kuramoto model [27] of synchronization have gained significant traction. These oscillator-based Ising machines (OIMs) [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38] based on Kuramoto model exploit the emergent collective behavior of synchronized oscillators to emulate spin interactions, with the system’s natural tendency towards stable, synchronized states corresponding to low-energy solutions of the Ising Hamiltonian. While Kuramoto-based OIMs have shown promise, oscillator models that incorporate amplitude dynamics in addition to phase [23, 39] have aimed to provide a more complete physical description of the underlying hardware and improved solution quality.
With both phase and amplitude dynamics, these models better reflect the physical behavior of real-world oscillators and offer greater flexibility in oscillator network design [40]. For clarity, we refer to models that integrate both amplitude and phase in the context of Ising machines as the Hopf oscillator model.
@@
  • Non-Monotonic Energy Decrease: Ideally, the system’s energy should decrease monotonically as it evolves towards a solution. However, we observed that even the real part, imaginary part, and absolute value of the previously defined complex-valued energy in the Hopf oscillator model do not consistently decrease during the optimization process. This behavior violates a fundamental principle of Ising machines and hinders both theoretical understanding and practical performance.
@@
  • Energy Redefinition: We introduce the complex conjugate of the oscillator representation variable $z$ to define a real-valued and physically meaningful energy function that directly corresponds to the objective function of the optimization problem.
@@
  • Generalization to Higher-Order Interactions: Our formulation naturally extends to include higher-order interactions among oscillators, thus broadening its applicability to more complex optimization problems.
@@ II.1 Ising Hamiltonian for Combinatorial Optimization
Later work revealed that finding the ground state of a spin glass is NP-hard [47]. Accordingly, a number of high-impact combinatorial problems can be mapped to an equivalent spin glass system [17, 48]. In particular, an ISG Hamiltonian can be trivially mapped to Quadratic Unconstrained Binary Optimization (QUBO) problems, a well-studied class within operations research [49].
@@ II.2 Solving 3-SAT problems with Ising Machine
A clause $c_{j}=(\ell_{j,1}\lor\ell_{j,2}\lor\ell_{j,3})$ is satisfied if at least one of its literals is true. Similar to individual literals, we can define a spin function $C_{j}(s)$ as
@@
Substituting Eq. (3), we can see that each $C_{j}$ is a spin polynomial of degree three
@@
From Eq. (5), when a clause is satisfied, its corresponding $C_{j}$ evaluates to zero, contributing nothing to the total energy. In contrast, if a clause is unsatisfied, $C_{j}$ contributes a positive value of $1$ to $\mathcal{H}$. Therefore, minimizing $\mathcal{H}$ effectively minimizes the number of unsatisfied clauses. The optimization problem can thus be formally stated as
@@
where $\mathbf{h}$, $\mathbf{J}$, and $\mathbf{P}$ are the coefficient vector, the coupling matrix, and the three-body interaction tensor, respectively. Minimizing $E$ corresponds to finding an assignment of spin variables $s_{i}$ that satisfies the maximum number of clauses. A solution with $E=0$ indicates that all clauses are satisfied, thereby providing a solution to the original 3-SAT problem.
In practice, analog Ising machines often represent binary spins $s$ by some physical quantities, thus relaxing the discrete phase space into a continuous one. We assume native device support for 3-body interactions and focus on the performance of the specific formulation. 3-body interactions can be implemented in analog electronic systems by means of multi-input gates (see Refs. [26, 24] for more details).
@@ II.3 Kuramoto Oscillator Model
where $\kappa$ and $J_{ij}$ describe pairwise and global interaction strengths respectively, and $\omega_{i}$ is the natural frequency of the $i^{th}$ oscillator. We set $\omega_{i}=0$ in Eq. (8) by assuming all oscillators have identical $\omega_{i}$ and viewing the system in a rotating reference frame. Using spin glass interaction coefficients for $J$ has led to oscillator-based accelerators for combinatorial optimization: oscillator Ising machines (OIM) [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 61, 62].
One can trivially derive the global Lyapunov function associated with Eq. (8) as
@@
The coupled oscillator system tends to minimize energy $E(\theta)$ over time [63], laying the foundation for OIMs that tackle combinatorial optimization problems. The original OIM architecture [28] utilized analog electronic LC-tank oscillators, and more recent proposals have broadened the choice of substrate to ring oscillators [64, 65], phase transition nano-oscillators [66, 67, 68], and magnetic tunnel junctions [69, 70].
@@ II.4 Sub-harmonic Injection Locking
To provide simple binarization, previous work [28] injects a sub-harmonic injection locking (SHIL) term to their spin dynamics,
$\frac{d\theta_{i}}{dt}=\kappa\sum_{i<j}^{N}J_{ij}\sin(\theta_{i}-\theta_{j})+\kappa_{S}(t)\overbrace{\sin(2\theta_{i})}^{\text{SHIL}}.$ (11)
@@ III.1 Hopf Oscillator Model
$\frac{dz_{i}}{dt}=f(z_{i})+g(z_{i}),$ (12)
where $z_{i}$ describes the phase and amplitude of the $i$-th oscillator, $f(z_{i})$ represents the local dynamics of the oscillator, and $g(z_{i})$ represents the coupling interaction between oscillators. In previous work [23] (which is also the baseline we compare with) , the local dynamics $f(z_{i})$ is defined as
@@
where the center frequency is set to $\omega_{i}=0$ for all oscillators. The parameters $\lambda_{i}$ and $\rho_{i}$ are chosen such that the system exhibits a stable limit cycle with unit amplitude. The term $g(z_{i})$ is derived from the model energy function $\mathcal{H}$. Previous work [23] directly maps the optimization problem to the system energy function $\mathcal{H}$ by replacing the Ising spin variable $s_{i}$ in Eq. (7) with the complex oscillator representation $z_{i}$ as follows:
@@
Although $\mathcal{H}$ can be interpreted as a function of oscillator states, several observations motivate us to investigate a revised formulation:
@@
    We derive the oscillator dynamics from the redefined energy and include the negative complex conjugate of the gradient of the energy function. This adjustment naturally drives the system toward the minima of the real-valued energy function.
Specifically, the energy function of our Hopf oscillator model is defined as follows:
@@
$+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{i}z_{j}^{*}z_{k}^{*})$
@@
Note that this construction is not unique; in Sec. IV.2 we discuss a general framework for generating appropriate energy functions. With the introduction of the complex conjugate, the energy function is no longer holomorphic. However, by applying Wirtinger calculus [41], we extend the concept of differentiation to the complex domain, allowing us to rigorously derive the system dynamics. We can treat the complex function $f(z)$ as a function of both $z$ and its conjugate $z^{*}$. For non-holomorphic functions, this requires differentiation with respect to both variables. The symmetry of our energy function with respect to $z$ and $z^{*}$ permits a simplification: we can compute the gradient by considering only the derivative with respect to either $z$ or $z^{*}$. Now we have,
$\frac{d\mathcal{H^{\prime}}}{dz_{i}}$ $=\frac{1}{2}h_{i}+\sum_{j=1}^{n}J_{ij}z_{j}^{*}$ (16)
$+\frac{1}{2}(\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{j}z_{k}^{*}+\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{j}z_{k}^{*}$
$+\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{j}^{*}z_{k}^{*})+\text{constant}.$
To drive the system towards minimizing the energy, we set
$\frac{dz_{i}}{dt}=-(\frac{d\mathcal{H^{\prime}}}{dz_{i}})^{*}$ (17)
and therefore obtain
[…]
with equality if and only if $\frac{d\mathcal{H^{\prime}}}{dz_{i}}=0$. That is, at a stationary point. Let $B_{r}(z^{\prime})$ be the ball of radius $r$ centered at $z^{\prime}$. If the stationary point $z^{\prime}$ is a local minimum, then the Hessian $\nabla^{2}\mathcal{H^{\prime}}(z^{\prime})$ is positive definite. From the definition of the Hessian in Wirtinger calculus [41] and the form of $\mathcal{H^{\prime}}$, we can conclude that $\nabla^{2}\mathcal{H^{\prime}}(z^{\prime})$ is continuous with respect to its arguments. Therefore, there exists an $\epsilon$ such that for all $y\in B_{\varepsilon}(z^{\prime})$, $\nabla^{2}\mathcal{H^{\prime}}(y)$ is also positive definite. Therefore, the dynamics (18) are locally Lyapunov stable within some neighborhood for all minima.
@@ III.2 Compatibility with Kuramoto model
The Kuramoto model was originally derived as a phase reduction for networks of weakly coupled oscillators [27]. Both the phase-only model (including the quadratic Kuramoto model) and the Hopf model only approximate real-world oscillators. However, the Hopf model is more expressive, as it accounts for amplitude variations. In the case of amplitude homogeneity, however, we expect that the Hopf model reduces to the phase-only model. Here, we show that the Hamiltonian of our model, Eq. (15) can be used to derive phase-only models found in previous literature in the context of cubic forms. In so doing, we demonstrate our model ① conforms to physical intuition and ② generalizes previous work.
@@
$+\frac{1}{2}(\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}|z_{i}||z_{j}||z_{k}|e^{i(\theta_{i}+\theta_{j}-\theta_{k})}$
@@
$+\sum_{i=1}^{n}\sum_{i=j}^{n}J_{ij}|z_{i}||z_{j}|\cos(\theta_{i}-\theta_{j})$
$+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}|z_{i}||z_{j}||z_{k}|\cos(\theta_{i}+\theta_{j}-\theta_{k})$
[…]
If we assume amplitude homogeneity ($|z_{i}|=|z_{j}|=|z_{k}|=1$), which will remove the amplitude dynamics from the model, then we obtain
$\mathcal{H^{\prime}}(z)=$ $\sum_{i=1}^{n}h_{i}\cos(\theta_{i})+\sum_{i=1}^{n}\sum_{i<j}^{n}J_{ij}\cos(\theta_{i}-\theta_{j})$ (21)
$+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}\cos(\theta_{i}+\theta_{j}-\theta_{k})$
[…]

### v2

# General Oscillator-Based Ising Machine Models with Phase-Amplitude Dynamics and Polynomial Interactions

@@ Abstract
We present an oscillator model with both phase and amplitude dynamics for oscillator-based Ising machines (OIMs). The model targets combinatorial optimization problems with polynomial cost functions of arbitrary order and addresses fundamental limitations of previous OIM models through a mathematically rigorous formulation with a well-defined energy function and corresponding dynamics. The model demonstrates monotonic energy decrease and reliable convergence to low-energy states. Empirical evaluations on 3-SAT problems show significant performance improvements over existing phase-amplitude models. Additionally, we propose a flexible, generalizable framework for designing higher-order oscillator interactions, from which we derive a practical method for oscillator binarization without compromising performance. This work strengthens both the theoretical foundation and practical applicability of oscillator-based Ising machines for complex optimization problems.
@@ I INTRODUCTION
The limitations of traditional von Neumann architectures in efficiently tackling combinatorial optimization problems have spurred the exploration of alternative computing paradigms. Among these, Ising machines, inspired by the Ising model from statistical mechanics, have emerged as a promising approach. The versatility of Ising machines extends across a diverse range of applications, including portfolio optimization [1, 2], drug discovery [3], vehicle routing [4, 5, 6], fraud detection [7], clustering [8], power grid optimization [9], wireless communication [10, 11], image restoration [12] and machine learning [13, 14, 15, 16, 17]. Additionally, Ising machines leverage multiple specialized hardware substrates, including quantum [18, 19], optical [20, 21, 22, 23, 24, 25, 26, 27, 28, 29], CMOS [30, 31, 32, 33, 34, 35, 36, 37] and other emerging hardware platforms [38, 39, 40, 41, 42, 43, 44]. By mapping the decision variables of an optimization problem onto interacting spins [45], Ising machines can efficiently search for solutions by allowing the system to naturally evolve towards its lowest energy state. This process corresponds to finding a configuration of spins that minimizes the system’s energy, which directly translates to a high-quality solution to the original optimization problem. This approach offers the potential to significantly accelerate the solution of complex optimization problems that are intractable for classical computers.
@@
Ising machines based on networks implementing the Kuramoto model [55] of synchronization have gained significant traction. These oscillator-based Ising machines (OIMs) [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] based on Kuramoto model exploit the emergent collective behavior of synchronized oscillators to emulate spin interactions, with the system’s natural tendency towards stable, synchronized states corresponding to low-energy solutions of the Ising Hamiltonian. While Kuramoto-based OIMs have shown promise, oscillator models that incorporate amplitude dynamics in addition to phase [51, 68, 69, 70, 71] have aimed to provide a more complete physical description of the underlying hardware. Research has revealed that phase-amplitude formulations for nanoelectromechanical (NEMS) oscillators can diverge significantly from simplified phase-only descriptions under specific operating regimes [68]. Moreover, incorporating amplitude dynamics into spintronic oscillator models (Slavin’s model [69]) yields a more faithful representation of coupled magnetization processes and the system demonstrates improved time complexity over the Goemans-Williamson algorithm [71]. For consistency with [51], we refer to models that integrate both amplitude and phase in the context of Ising machines as the Hopf oscillator model.
@@
  • Non-Monotonic Energy Decrease: Ideally, the system’s energy should be able to decrease monotonically in the absence of introduced noise. In prior model [51], we observed that neither individual parts (real or imaginary) nor absolute value of the energy consistently decrease during the optimization process.
@@
  • Energy Definition: We introduce the complex conjugate of the oscillator representation variable $z$ to define a real-valued and physically meaningful energy function that directly corresponds to the objective function of the optimization problem.
@@
  • Generalization to Higher-Order Interactions: Our formulation naturally extends higher-order interactions among oscillators, thus broadening its applicability to more complex optimization problems.
@@ II.1 Ising Hamiltonian for Combinatorial Optimization
Later work revealed that finding the ground state of a spin glass is NP-hard [79, 80]. Accordingly, a number of high-impact combinatorial problems can be mapped to an equivalent spin glass system [45, 81]. In particular, an ISG Hamiltonian can be trivially mapped to Quadratic Unconstrained Binary Optimization (QUBO) problems, a well-studied class within operations research [82].
@@ II.2 Solving 3-SAT problems with Ising Machine
A clause $c_{j}=(\ell_{j,1}\lor\ell_{j,2}\lor\ell_{j,3})$ is satisfied if at least one of its literals is true. Similarly to individual literals, we can define a spin function $C_{j}(s)$ as
@@
Substituting Eq. 3, we can see that each $C_{j}$ is a spin polynomial of degree three
@@
From Eq. 5, when a clause is satisfied, its corresponding $C_{j}$ evaluates to zero, contributing nothing to the total energy. In contrast, if a clause is unsatisfied, $C_{j}$ contributes a positive value of $1$ to $\mathcal{H}$. Therefore, minimizing $\mathcal{H}$ effectively minimizes the number of unsatisfied clauses. The optimization problem can thus be formally stated as
@@
where h{h}, J{J}, and P{P} are the coefficient vector, the coupling matrix, and the three-body interaction tensor, respectively. Minimizing $E$ corresponds to finding an assignment of spin variables $s_{i}$ that satisfies the maximum number of clauses. A solution with $E=0$ indicates that all clauses are satisfied, thereby providing a solution to the original 3-SAT problem. Appendix A provides a concrete example by mapping a single 3-SAT clause to the PUBO Hamiltonian.
In practice, analog Ising machines often represent binary spins $s$ by some physical quantities, thus relaxing the discrete phase space into a continuous one. We assume native device support for 3-body interactions and focus on the performance of the specific formulation. 3-body interactions can be implemented in analog electronic systems by means of multi-input gates (see [54, 52] for more details).
@@ II.3 Kuramoto Oscillator Model
where $\kappa$ and $J_{ij}$ describe global and pairwise interaction strengths respectively, and $\omega_{i}$ is the natural frequency of the $i^{th}$ oscillator. We set $\omega_{i}=0$ in Eq. 8 assuming that all oscillators have identical $\omega_{i}$ and viewing the system in a rotating reference frame. Using spin glass interaction coefficients for $J$ has led to oscillator-based accelerators for combinatorial optimization: oscillator Ising machines (OIM) [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 93, 94].
We can trivially derive the global Lyapunov function associated with Eq. 8 as
@@
The coupled oscillator system tends to minimize energy $E(\theta)$ over time [95], laying the foundation for OIMs that tackle combinatorial optimization problems. The original OIM architecture [56] utilized electronic LC-tank oscillators, and more recent proposals have broadened the choice of substrate to ring oscillators [96, 97], phase transition nano-oscillators [98, 99, 100], and magnetic tunnel junctions [101, 102].
@@ II.4 Sub-harmonic Injection Locking
To provide simple binarization, previous work [56] injects a sub-harmonic injection locking (SHIL) term into their spin dynamics,
$\frac{d\theta_{i}}{dt}=\kappa\sum_{i<j}^{N}J_{ij}\sin(\theta_{i}-\theta_{j})-\kappa_{S}(t)\overbrace{\sin(2\theta_{i})}^{\text{SHIL}}.$ (11)
@@ III.1 Hopf Oscillator Model
$\frac{dz_{i}}{dt}=f(z_{i})+g(z),$ (12)
where $z_{i}$ describes the phase and amplitude of the $i$-th oscillator, $f(z_{i})$ represents the local dynamics of the oscillator, and $g(z)$ represents the coupling interaction between oscillators. In previous work [51] (which is also the baseline we compare with), the local dynamics $f(z_{i})$ is defined as
@@
where the center frequency is set to $\omega_{i}=0$ for all oscillators. The parameters $\lambda_{i}$ and $\rho_{i}$ are chosen such that the system exhibits a stable limit cycle with unit amplitude. The term $g(z)$ is derived from the model energy function $\mathcal{H}$. Previous work [51] directly maps the optimization problem to the system energy function $\mathcal{H}$ by replacing the Ising spin variable $s_{i}$ in Eq. 7 with the complex oscillator representation $z_{i}$ as follows:
@@
Several observations motivate us to investigate a new formulation:
@@
    Using the new energy function, we derive system dynamics using gradient of the energy function. The resulting gradient system naturally drives the state toward minima of the real-valued energy function.
Specifically, one of the energy functions of our Hopf oscillator model is defined as follows:
@@
$+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{i}z_{j}^{*}z_{k}^{*})+\text{constant}.$
Note that this construction is not unique; in Sec. IV.2 we discuss a general framework for generating appropriate energy functions.
With the introduction of the complex conjugate, the energy function is no longer holomorphic. However, by applying Wirtinger calculus [73], we can rigorously derive the system dynamics. The core idea of Wirtinger calculus is to treat the complex function $f(z)$ as a function of both $z$ and its conjugate $z^{*}$ as independent variables. For non-holomorphic functions, this requires differentiation with respect to both variables. Appendix B contains a brief overview of Wirtinger calculus.
The symmetry of our energy functions with respect to $z$ and $z^{*}$ permits a simplification: we can compute the gradient by considering only the derivative with respect to either $z$ or $z^{*}$. Now we have,
$\frac{\partial\mathcal{H^{\prime}}}{\partial z_{i}}$ $=\frac{1}{2}h_{i}+\sum_{j=1}^{n}J_{ij}z_{j}^{*}$ (16)
$+\frac{1}{2}\left(\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{j}z_{k}^{*}+\sum_{i=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{i}z_{k}^{*}\right$.
$+\left.\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{j}^{*}z_{k}^{*}\right).$
For convenience, the coefficient $P$ is constructed to be symmetric under all permutations of its indices. It follows that $\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}\,z_{j}\,z_{k}^{*}=\sum_{i=1}^{n}\sum_{k=1}^{n}P_{ijk}\,z_{i}\,z_{k}^{*}$. Therefore, summing over any two of the three indices yields the same result. To drive the system towards minimizing the energy, we set
$\frac{dz_{i}}{dt}=-\left(\frac{\partial\mathcal{H^{\prime}}}{\partial z_{i}}\right)^{*}.$ (17)
Furthermore, we have the identity
$\frac{dz_{i}}{dt}=\left(\frac{dz_{i}^{*}}{dt}\right)^{*},$ (18)
which follows from the definition of $z$ and $z^{*}$. Additionally, we have
$\left(\frac{\partial\mathcal{H^{\prime}}}{\partial z_{i}}\right)^{*}=\partialderivative{\mathcal{H^{\prime}}}{z_{i}^{*}},$ (19)
which follows from the conjugation property of Wirtinger derivatives (see Appendix B). Combining these results with Eq. 17, we obtain the total derivative:
$\frac{d\mathcal{H^{\prime}}}{dt}$ $=\sum_{i}\frac{\partial\mathcal{H^{\prime}}}{\partial z_{i}}\frac{dz_{i}}{dt}+\sum_{i}\frac{\partial\mathcal{H^{\prime}}}{\partial z_{i}^{*}}\frac{dz_{i}^{*}}{dt}$ (20)
$=-2\sum_{i}\frac{\partial\mathcal{H^{\prime}}}{\partial z_{i}}\left(\frac{\partial\mathcal{H^{\prime}}}{\partial z_{i}}\right)^{*}$
@@
with equality if and only if $\frac{d\mathcal{H^{\prime}}}{dz_{i}}=0$, i.e., at a stationary point. Since the Hessian operator remains continuous in Wirtinger calculus [73], the dynamics (Eq. 20) are locally Lyapunov stable within some neighborhood for all minima. Therefore, Eq. 15 is a suitable Lyapunov function for complex-domain Ising machines. In contrast, the energy formulation proposed in previous work [51] is complex-valued and cannot serve as a viable Lyapunov function [^1] Disregarding the physical interpretation of the complex-valued $\mathcal{H}$, we examine the behavior of the baseline model [51] on a simple 3-variable problem, $E=s_{1}s_{2}s_{3}$, where $s_{i}\in\{-1,+1\}$. Both real and imaginary components exhibit oscillatory behavior, which leads to unstable dynamics. This can be demonstrated mathematically and empirically, and the observation extends to general cases..
@@ III.2 Compatibility with Kuramoto model
The Kuramoto model was originally derived as a phase reduction for networks of weakly coupled oscillators [55]. Both the phase-only model (including the quadratic Kuramoto model) and the Hopf model only approximate real-world oscillators. However, the Hopf model is more expressive, as it accounts for amplitude variations. One would expect that in the case of amplitude homogeneity, the Hopf model reduces to the phase-only model. Here, we show that indeed the Hamiltonian of our model, Eq. 15 can be used to derive phase-only models found in previous literature in the context of cubic forms. In so doing, we demonstrate our model ① conforms to physical intuition and ② generalizes previous work.
@@
$+\frac{1}{2}(\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}|z_{i}||z_{j}||z_{k}|e^{i(\theta_{i}+\theta_{j}-\theta_{k})}$
@@
$+\sum_{i=1}^{n}\sum_{j=1}^{n}J_{ij}|z_{i}||z_{j}|\cos(\theta_{i}-\theta_{j})$
$+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}|z_{i}||z_{j}||z_{k}|\cos(\theta_{i}+\theta_{j}-\theta_{k})$
$+\text{constant}.$
If we further assume amplitude homogeneity ($|z_{i}|=|z_{j}|=|z_{k}|=1$), which will remove the amplitude dynamics from the model, then we obtain
$\mathcal{H^{\prime}}(z)$ $=\sum_{i=1}^{n}h_{i}\cos(\theta_{i})+\sum_{i=1}^{n}\sum_{j=1}^{n}J_{ij}\cos(\theta_{i}-\theta_{j})$ (23)
$+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}\cos(\theta_{i}+\theta_{j}-\theta_{k})+\text{constant},$